IV. Calculation and Analysis.(30 points) The directors of Sunshine Theme Parks Limited are considering opening a new park in the holiday resort. They have lease land license to operate the theme park for five years. Additional information theme for five years, and have been granted a Forecasts show that the initial cost of the investment will be $6 000 000. The investment of $6 000 000 includes 6 rides, costing $150 000 each. The rides are to be depreciated over 5 years using the straight-line depreciation method at a rate of 20% per year. Running expenses, including depreciation for the rides, for the first three years are forecast to be $1 000 000 per year. In year four, running expenses, including depreciation, are forecast to rise by $200 000. In year five, running expenses are forecast to stay at the same level as in year four. 第8页共10页
浙江科技学院考试 For the first three years it is forecast that the theme park will operale for 210 days per year, with an average of 500 visitors per day, each paying a $2 5 entrance fee. In year four, it is forecast that the theme park will operate for 210 days, with an average of 600 visitors a day, each paying a S30 entrance foe. These figures are expected to stay the same for year five. .It is company policy to have a payback period of five years on investment projects. And the opportunity cost of capital is 16%. Required: (1) Calculate the payback period of this project (2) Calculate the net present value of this project. (3) Evaluate the project on behalf of Sunshine Theme Parks Limited. (4) Which factors should you consider, when making the investment decisions?
Answer #1
Years 0 1 2 3 4 5
Initial Investment (A) -6000000
Revenue (B) 2625000 2625000 2625000 3780000 3780000
Depreciation Expenses ( C ) 180000 180000 180000 180000 180000
Other Running Expenses (D) 820000 820000 820000 1020000 1020000
EBIT (E=B-C-D) 1625000 1625000 1625000 2580000 2580000
Tax (Tax rate not given) 0 0 0 0 0
Earning After Tax (F) 1625000 1625000 1625000 2580000 2580000
Operating Income (G=F+C) 1805000 1805000 1805000 2760000 2760000
PVIF @ 16% 1 0.862069 0.743163 0.640658 0.552291 0.476113
PV -6000000 1556034 1341409 1156387 1524323 1314072
NPV 892226
Cumulative Present Value -6000000 -4443966 -3102556 -1946169 -421846 892226

1)

Payback Period = 4years + 421846/1314072
PBP = 4 + 0.32102 year
PBP = 4.32102 year

2) NPV = 892226 (as calculated in the table above)

3) As per the policy of sunshine theme park ltd. a project is accepted if the payback period of the project is less than 5 years. In the given case the PBP of the project is 4.32 years, thus the project is viable.

4) Following factors should be considered while making an investment decision

a) Timing of Cash Flows : Cash Flows received during the initial period of a project is more valuable then those received in the later periods as the time value of money comes into effect.

b) Net present value of the project : A company should accept those projects which has positive NPV.

c) Payback Period of the project : PBP gives an indication about in how much time a company will recover its initial outlay in terms of present value of money.

d) Internal Rate of Return for the project : A company should accept a project when the project IRR is greater than the required rate of return i.e. opportunity cost of capital.